# Modular Periodicity of Random Initialized Recurrences

### v1

@@ Abstract
Classical studies of the Fibonacci sequence focus on its periodicity modulo $m$ (the Pisano periods) with canonical initialization. We investigate instead the complete periodic structure arising from all $m^{2}$ possible initializations in $\mathbb{Z}/m\mathbb{Z}$. We discover perfect chiral symmetry between the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$, and observe fractal self-similarity in the extension from prime to prime power moduli. Additionally, we classify prime moduli due to their periodicity, and show that they exhibit weight preservation under modular extension. Furthermore, we define a minima distribution $P(n)$ governed by Lucas ratios, which satisfies the symmetric relation P(n)=P(1n)P(n)=P(1-n). For cyclotomic recurrences, we conjecture explicit counting functions for the number of distinct periods, with connections to necklace enumeration. These findings suggest potential links to Viswanath’s random recurrence, modular forms and L-functions.
Keywords— Fibonacci sequence, Pisano period, cyclotomic polynomial, chiral symmetry, recurrence relation, modular arithmetic
@@ 1 Introduction
However, the classical Pisano problem addresses only a single sequence with fixed initialization. In $\mathbb{Z}/m\mathbb{Z}$, there are $m^{2}$ possible initial conditions — and in general there are $m^{k}$ initial conditions for a recurrence of order $k$ — clustering into a variety of irreducible and cyclically equivalent sequences. For example, the period [011011] reduces to [011], which is cyclic equivalent to [101] and [110], capturing the three initial conditions (0,1)(0,1),$(1,1)$ and $(1,0)$.
This naturally leads to a broader question: what is the complete structure of all periodic sequences arising from a given recurrence modulo $m$? How many distinct periods exist and what are their lengths? Understanding this "period landscape" reveals algebraic structure invisible when studying only the canonical Fibonacci sequence.
To approach this systematically, we focus on recurrences with restricted coefficients, namely monic quadratic polynomials with coefficients in $\{-1,+1\}$ (also known as monic Littlewood polynomials [3]). Two of these four polynomials are $x^{2}-x-1$ and $x^{2}+x-1$, which generate the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$. Notably, these are the two deterministic branches of Viswanath’s random recurrence $a_{n}=a_{n-1}\pm a_{n-2}$, where the sign is chosen at random at each step, and where the sequence grows exponentially with rate approximately $1.13198824$ [4]. The other two polynomials are the cyclotomic polynomials $\Phi_{3}(x)=x^{2}+x+1$ and $\Phi_{6}(x)=x^{2}-x+1$, whose roots are primitive roots of unity. More generally, the $n$th cyclotomic polynomial is defined as
@@
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences. For the cyclotomic cases, we conjecture explicit formulas for the number of distinct periods modulo $m$ (1,4,7 and 9), verified through extensive computation. Between the Fibonacci and its parity recurrence, we discover an unexpected chiral symmetry: despite having different characteristic roots, they produce identical period structures for all tested moduli. We conjecture that the periodicity of prime moduli is due to the number of zeros in the Pisano period, together with a mysterious variable $\alpha$ (11), and display the fractal-like self-similarity for prime power moduli (16). Furthermore, we conjecture that these periods exhibit weight preservation (18) when extending from divisor moduli to composite moduli, and that both recurrences share an identical minima distribution (21). These findings suggest deep connections between deterministic periodicity in finite rings and questions of exponential growth in random recurrences.
@@ 2 Periodicity of Cyclotomic Recurrences
This section addresses the periodicity of recurrences with characteristic polynomial being cyclotomic. For example, following equations 1 and 2, the corresponding recurrence for the first cyclotomic polynomial is $a_{n}=a_{n-1}$, for the third it is $a_{n}=-a_{n-2}-a_{n-1}$, and for the ninth it is $a_{n}=-a_{n-6}-a_{n-3}$. Notably the first cyclotomic polynomial having other coefficients than 1,0, or -1 is $\Phi_{105=3\cdot 5\cdot 7}$.
@@
$\#_{\pi,10}(10)$ $=\frac{10^{4}-2^{4}-5+1}{10}+1+\frac{4}{2}+\frac{2^{4}-1}{5}=\frac{9980}{10}+6=998+6=1004$ (7)
@@
$\#_{\pi,9}(12)$ $=\frac{12^{6}-(1\cdot 3+3\cdot 8)}{9}+(1\cdot 3+3\cdot 8)=\frac{2985957}{9}+27=331773+27=331800$ (10)
@@
$\#^{6}_{\pi}(4)$ $=\frac{4^{6}-(1\cdot 4+2\cdot 6+3\cdot 20)}{6}+(1\cdot 4+2\cdot 6+3\cdot 20)=\frac{4020}{6}+76=670+76=746$ (12)
@@ 3 Periodicity of Fibonacci Recurrences
We now turn our focus to the Fibonacci recurrence $a_{n}=a_{n-2}+a_{n-1}$ and its parity transform $a_{n}=a_{n-2}-a_{n-1}$ with random integer initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$. The implied mirror-symmetry can be seen in Table 1 for all moduli $m$, where the number of periods and lengths are similar for both recurrences, and where every period corresponds to a chiral period from the other recurrence, respectively.
@@
Conjecture 11 (Period Count for Prime Fibonacci Recurrences).
Let $p$ be a prime and let #(p)\#(p) denote the number of distinct periods when the Fibonacci recurrence or its parity transform is computed modulo $p$ over all possible initial conditions $(a_{0},a_{1})\in(\mathbb{Z}/p\mathbb{Z})^{2}$. Primes divide into two types:
Type 1 (Two period lengths): The Pisano period satisfies $\pi(p)=\frac{2(p+1)}{\alpha}$ where $\alpha$ is an odd positive integer. The number of distinct periods is:
$(\alpha,p)=\frac{(p^{2}-1)\alpha}{2(p+1)}+1$ (13)
Additionally, for odd and even $\alpha$, the Pisano period satisfies $\pi(p)=\frac{p-1}{\alpha}$ and:
$(\alpha,p)=\frac{(p^{2}-1)\alpha}{p-1}+1$ (14)
Type 2 (Three period lengths): The Pisano period satisfies $\pi(p)=\frac{p-1}{\alpha}$ where $\alpha$ is a positive integer. The number of distinct periods is:
$_{2}(\alpha,p)=p\cdot\alpha+2\alpha+1$ (15)
Type-2 primes are precisely those whose Pisano period contains exactly one zero, corresponding to OEIS sequence A053032 (seperately excluding $p=2$ but including $p=5$).
Remark 12 (Connection to Quadratic Residues).
The classification is intimately connected to the Legendre symbol $\left(\frac{5}{p}\right)$. Type-1 primes with odd $\alpha$ (satisfying $\pi(p)=\frac{2(p+1)}{\alpha}$) correspond precisely to primes where 5 is a quadratic non-residue, while Type-2 primes and Type-1 primes with even $\alpha$ (both satisfying $\pi(p)=\frac{p-1}{\alpha}$) correspond to primes where 5 is a quadratic residue. This follows from the classical result that $\pi(p)\mid(p-1)$ when $\left(\frac{5}{p}\right)=1$ and $\pi(p)\mid 2(p+1)$ when $\left(\frac{5}{p}\right)=-1$.
A finer classification emerges via congruences modulo 20. Primes satisfying $p\equiv 3,7,13,17\pmod{20}$ are Type-1 with $\pi(p)=\frac{2(p+1)}{\alpha}$. Among primes with $\left(\frac{5}{p}\right)=1$, those satisfying $p\equiv 11,19\pmod{20}$ are predominantly Type-2, while $p\equiv 1,9\pmod{20}$ exhibit mixed behavior. The precise determination of $\alpha$ for a given prime remains an open problem.
Remark 13.
For Type-1 primes, the period structure consists of the zero vector (length 1) and #1\#-1 periods of length $\pi(p)$. For Type-2 primes, there are three distinct period lengths: the zero vector (length 1) appears once, length $\frac{\pi(p)}{2}$ appears exactly $2\alpha$ times, and length $\pi(p)$ appears $p\cdot\alpha$ times.
Example 14 (Type-1 Prime: $p=89$, $n=2$).
The Pisano period is $\pi(89)=\frac{89-1}{2}=44$. The number of distinct periods is:
$(2,89)=\frac{(89^{2}-1)\cdot 2}{89-1}+1=\frac{15840}{88}+1=180+1=181$ (16)
@@
Example 15 (Type-2 Prime: $p=11$, $n=1$).
The Pisano period is $\pi(11)=\frac{11-1}{1}=10$. The number of distinct periods is:
$(1,11)=11\cdot 1+2\cdot 1+1=14$ (17)
The period structure consists of one period of length $1$ (zero vector), $2$ periods of length $5$, and $11$ periods of length $10$.
@@
The period structure for the Fibonacci recurrence and its parity transform exhibit hierarchical self-similarity between prime and prime power moduli. At each transition $p^{k}\to p^{k+1}$, all existing periods are preserved, and each period of length $\ell>1$ at $p^{k}$ generates new periods of length $p\ell$ at $p^{k+1}$, with multiplicities scaling by factor $p$. For Type-2 primes, however, the multiplicity for the middle period remains constant at exactly $2\alpha$ across all powers $p^{k}$.
Example 17 (Prime Power Extension: $p=19$, $\alpha=1$).
@@
Consider the Fibonacci recurrence $a_{n}=a_{n-2}+a_{n-1}$ and its chiral recurrence $a_{n}=a_{n-2}-a_{n-1}$ with random integer initialization $\{a_{0},a_{1}\}\in\mathbb{Z}^{2}$. Although these sequences diverge as $n\to\pm\infty$, they possess well-defined absolute minima. Let $F_{n}$ and $L_{n}$ denote the $n$-th Fibonacci and Lucas numbers, respectively. Define $P(n)$ as the probability that a randomly initialized sequence has its absolute minimum at position $n$. Then:
@@ 4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect chiral symmetry between the Fibonacci and parity recurrences. We classified primes into two types based on their period structure, observed fractal self-similarity at prime powers, established weight preservation under modular extension, and derived probability distributions for sequence minima governed by Lucas ratios.
A central mystery remains: the classification parameter $\alpha$ that determines a prime’s type and governs the exact period count. We observe that the type classification is intimately connected to whether 5 is a quadratic residue modulo $p$, however, the precise value of $\alpha$ for a given prime remains unknown. Understanding whether $\alpha$ relates to higher-order residue properties, the structure underlying Viswanath’s random recurrence, or other arithmetic invariants could reveal deeper connections between Fibonacci periodicity and algebraic number theory.
@@
The framework extends naturally beyond our quadratic cases. While we investigated the two Littlewood polynomials $x^{2}\pm x-1$ among quadratics, higher-degree polynomials with restricted coefficients also reveal rich periodicity structures. The most significant example are the order-6 recurrences $a_{n}=a_{n-6}+a_{n-3}$ and $a_{n}=a_{n-6}-a_{n-3}$, which are the negations of the cyclotomic recurrences from $\Phi_{9}$ and $\Phi_{18}$ (analogous to how Fibonacci and Parity arise as negations of $\Phi_{3}$ and $\Phi_{6}$). These exhibit identical periodicity for all tested moduli $m\leq 19$, demonstrating that chiral symmetry might extend beyond the quadratic case. Remarkably, classical Fibonacci periods (such as $01120221$ mod 3 and $1342$ and $01123033140443202241$ mod 5) appear within these sequences, suggesting hierarchical connections. However, fully understanding their period structure requires complex computations. Certain primes exhibit five distinct period lengths, rather than the two or three observed in the quadratic case. Additionally, self-similarity at prime powers follows more subtle rules than the direct scaling observed for Fibonacci-Parity. Whether weight preservation and generalized classification formulas extend to this setting, and whether the negations of the cyclotomic recurrences $\Phi_{3^{j}}$ and $\Phi_{2\cdot 3^{j}}$, for all $j\geq 1$, yield chirality remains important open questions. A rigorous understanding may require companion matrix theory and conjugation properties under sign transformations.
Another interesting family is that of the Fekete polynomials, which are related to cyclotomic polynomials and L-functions, as demonstrated in [7]. Studying the complete period landscapes of these and other polynomial families could reveal universal principles governing recurrence periodicity in finite rings.
Finally, preliminary computational evidence suggests intriguing relationships between Fibonacci period lengths and cyclotomic recurrences. Certain Fibonacci periods appear systematically across families of cyclotomic polynomials—for instance, the Pisano period 011 only appears modulo 2 in all $\Phi_{n}$ with $3\mid n$, and modulo 3 in all $\Phi_{3^{j}}$ for all $j>1$. In contrast, periods from primes in OEIS A053032 (Type-2 primes where the Pisano period contains exactly one zero) rarely or never appear in cyclotomic landscapes. Whether this pattern reflects a deeper structural link between exponentially growing sequences (Fibonacci) and unit-root periodicity (cyclotomic) remains an open question worthy of further investigation.

### v2

@@ Abstract
Classical studies of the Fibonacci sequence focus on its periodicity modulo $m$ (the Pisano periods) with canonical initialization. We investigate instead the complete periodic structure arising from all $m^{2}$ possible initializations in $\mathbb{Z}/m\mathbb{Z}$. We discover perfect chiral symmetry between the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ and observe fractal self-similarity in the extension from prime to prime power moduli. Additionally, we classify prime moduli based on their quadratic reciprocity and demonstrate that periodic sequences exhibit weight preservation under modular extension. Furthermore, we define a minima distribution $P(n)$ governed by Lucas ratios, which satisfies the symmetric relation P(n)=P(1n)P(n)=P(1-n). For cyclotomic recurrences, we propose explicit counting functions for the number of distinct periods with connections to necklace enumeration. These findings imply potential connections to Viswanath’s random recurrence, modular forms and L-functions.
Keywords— Fibonacci sequence, Pisano period, cyclotomic polynomial, chiral symmetry, random recurrence, modular arithmetic
@@ 1 Introduction
However, the classical Pisano problem addresses only a single sequence with fixed initialization. In $\mathbb{Z}/m\mathbb{Z}$, there are $m^{2}$ possible initial conditions — and in general there are $m^{k}$ initial conditions for a recurrence of order $k$ — clustering into a variety of irreducible and cyclically equivalent sequences. For example, the period [011011] reduces to [011], which is cyclic equivalent to [101] and [110], capturing the three initial conditions (0,1)(0,1),$(1,1)$ and $(1,0)$. This naturally leads to a broader question: what is the complete structure of all periodic sequences arising from a given recurrence modulo $m$? How many distinct periods exist and what are their lengths? Understanding this "period landscape" reveals algebraic structure invisible when studying only the canonical Fibonacci sequence.
To approach this systematically, we focus on recurrences with restricted coefficients, namely monic quadratic polynomials with coefficients in $\{-1,+1\}$ (also known as Littlewood polynomials [3]). Two of these four polynomials are $x^{2}-x-1$ and $x^{2}+x-1$, which generate the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$. The other two are the cyclotomic polynomials $\Phi_{3}(x)=x^{2}+x+1$ and $\Phi_{6}(x)=x^{2}-x+1$, whose roots are primitive roots of unity. Interestingly, the Fibonacci recurrence (which is the negation from $\Phi_{3}$) and the recurrence $a_{n}=a_{n-1}-a_{n-2}$ (from $\Phi_{6}$) form the two deterministic branches of Viswanath’s random recurrence $a_{n}=a_{n-1}\pm a_{n-2}$, which exhibits exponential growth with rate approximately $1.13198824$ [4]. This connection between non-cyclotomic and cyclotomic recurrences motivates our parallel investigation of both families. More generally, the $n$th cyclotomic polynomial is defined as
@@
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences across three main sections. In section 2, we conjecture explicit counting functions for the number of distinct periods modulo $m$ for various cyclotomic families, including $\Phi_{p}$ (1), $\Phi_{2p}$ (4), $\Phi_{p^{j}}$ (7) and $x^{n}-1$ (9), all supported by extensive computational verification. In section 3, the non-cyclotomic Fibonacci and parity recurrences are explored, revealing unexpected chiral symmetry. Despite having algebraically distinct characteristic roots, these recurrences produce identical period structures for all tested moduli. We classify prime moduli based on their quadratic reciprocity and an additional parameter $\alpha$ (11), observe fractal self-similarity at prime power moduli (15), and conjecture weight preservation under modular extension (17). Furthermore, we establish that both recurrences share an identical minimum distribution governed by Lucas number ratios (20). In section 4, we discuss broader implications and open questions, and suggest that deterministic periodicity in finite rings has potential relations to stochastic growth in random recurrences and to classical number-theoretical objects such as modular forms and L-functions.
@@ 2 Periodicity of Cyclotomic Recurrences
This section addresses the periodicity of recurrences with characteristic polynomial being cyclotomic. For example, following equations 1 and 2, the corresponding recurrence for the first cyclotomic polynomial is $a_{n}=a_{n-1}$, for the third it is $a_{n}=-a_{n-1}-a_{n-2}$ and for the ninth it is $a_{n}=-a_{n-3}-a_{n-6}$. Notably the first cyclotomic polynomial having other coefficients than 1,0, or -1 is $\Phi_{105=3\cdot 5\cdot 7}$.
@@
$\#(\Phi_{10},10)$ $=\frac{10^{4}-2^{4}-5+1}{10}+1+\frac{4}{2}+\frac{2^{4}-1}{5}=\frac{9980}{10}+6=998+6=1004$ (7)
@@
$\#(\Phi_{9},12)$ $=\frac{12^{6}-(1\cdot 3+3\cdot 8)}{9}+(3+8)=\frac{2985957}{9}+11=331773+27=331784$ (10)
@@
$\#(6,4)$ $=\frac{4^{6}-(1\cdot 4+2\cdot 6+3\cdot 20)}{6}+(4+6+20)=\frac{4020}{6}+30=670+30=700$ (12)
@@ 3 Periodicity of Fibonacci Recurrences
We now turn our focus to the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$. The implied mirror-symmetry can be seen in Table 1 for all moduli $m$, where the number of periods and lengths are similar for both recurrences, and where every period corresponds to a chiral period from the other recurrence.
@@
Conjecture 11 (Period Count for Fibonacci Recurrences modulo p).
Let $p$ be a prime and let #(p)\#(p) denote the number of distinct periods when the Fibonacci recurrence or its parity transform is computed modulo $p$ over all possible initial conditions $(a_{0},a_{1})\in(\mathbb{Z}/p\mathbb{Z})^{2}$. The count is determined by the Legendre symbol $\left(\frac{5}{p}\right)$ and a positive integer $\alpha$ governing the Pisano period:
[…]
$_{A}(p)=\frac{\alpha}{2}(p-1)+1$ (13)
These primes exhibit two period lengths: the zero vector and non-trivial periods of length $\pi_{A}(p)$.
[…]
Subclass B1 (two period lengths): The number of distinct periods is:
$_{B1}(p)=\alpha(p+1)+1$ (14)
These primes exhibit two periodic lengths: the zero vector and non-trivial periods of length $\pi_{B}(p)$.
Subclass B2 (three period lengths): The Pisano period contains exactly one zero (OEIS A053032 for $p\geq 11$). The number of distinct periods is:
$_{B2}(p)=\alpha(p+2)+1$ (15)
These primes exhibit three period lengths: the zero vector, an intermediate length $\frac{\pi_{B}(p)}{2}$ appearing $2\alpha$ times, and the length $\pi_{B}(p)$ appearing $p\alpha$ times.
All primes satisfying $p\equiv 11,19\pmod{20}$ belong to subclass B2, while those satisfying $p\equiv 1,9\pmod{20}$ can belong to either subclass.
Example 12 (Class A Prime: $p=47$, $\alpha=3$).
The Pisano period is $\pi_{A}(47)=\frac{2(47+1)}{3}=32$. The number of distinct periods is:
$_{A}(47)=\frac{3}{2}(47-1)+1=69+1=70$ (16)
The period structure consists of the zero vector and $69$ periods of length $32$.
Example 13 (Class B1 Prime: $p=89$, $\alpha=2$).
The Pisano period is $\pi_{B}(89)=\frac{89-1}{2}=44$. The number of distinct periods is:
$_{B1}(89)=2\cdot(89+1)+1=180+1=181$ (17)
@@
Example 14 (Class B2 Prime: $p=11$, $\alpha=1$).
The Pisano period is $\pi_{B}(11)=\frac{11-1}{1}=10$. The number of distinct periods is:
$_{B2}(11)=1\cdot(11+2)+1=13+1=14$ (18)
The period structure consists of the zero vector, $2$ periods of length $5$ and $11$ periods of length $10$.
@@
The period structure for the Fibonacci recurrence and its parity transform exhibit hierarchical self-similarity between prime and prime power moduli. At each transition $p^{k}\to p^{k+1}$, all existing periods are preserved and each period of length $\ell>1$ at $p^{k}$ generates new periods of length $p\ell$ at $p^{k+1}$, with multiplicities scaling by factor $p$. For class B2 primes, however, the multiplicity for the middle period remains constant at exactly $2\alpha$ across all powers $p^{k}$.
Example 16 (Prime Power Extension for B2 prime $p=19$, $\alpha=1$).
@@
Consider the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its chiral recurrence $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $\{a_{0},a_{1}\}\in\mathbb{Z}^{2}$. Although these sequences diverge as $n\to\pm\infty$, they possess well-defined absolute minima. Let $F_{n}$ and $L_{n}$ denote the $n$-th Fibonacci and Lucas numbers respectively. Define $P(n)$ as the probability that a randomly initialized sequence has its absolute minimum at position $n$. Then:
@@ 4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect chiral symmetry between the Fibonacci and parity recurrences. We classified prime moduli based on their quadratic reciprocity, observed fractal self-similarity at prime power moduli, established weight preservation under modular extension and derived probability distributions for sequence minima governed by Lucas ratios.
@@
The framework of random initialization naturally extends beyond our quadratic cases. While we investigated the two Littlewood polynomials $x^{2}\pm x-1$ among quadratics, higher-degree polynomials with restricted coefficients also reveal rich periodicity structures. The most significant examples are the order-6 recurrences $a_{n}=a_{n-3}+a_{n-6}$ and $a_{n}=-a_{n-3}+a_{n-6}$ (which arise as negations from $\Phi_{9}$ and $\Phi_{18}$). Analogous to how the Fibonacci and parity recurrences are negations from $\Phi_{3}$ and $\Phi_{6}$, these order-6 recurrences also exhibit identical periodicity for all tested moduli $m\leq 19$, demonstrating that chiral symmetry might extend beyond the quadratic case. A rigorous proof of such symmetries may require companion matrix theory and conjugation properties under sign transformations. Observations show that classical Fibonacci periods (such as $01120221$ mod 3 and $1342$ and $01123033140443202241$ mod 5) appear within these order-6 recurrences, suggesting hierarchical connections. However, their period structure is more complex: certain primes exhibit five distinct period lengths rather than the two or three observed in the quadratic case, and self-similarity at prime powers follows subtler rules. Whether weight preservation and generalized classification formulas extend to this setting remains an open question. Moreover, both order-6 recurrences can be viewed as three interleaved Fibonacci or parity recurrences respectively (via initial ratios $a_{0}/a_{3}$, $a_{1}/a_{4}$ and $a_{2}/a_{5}$), suggesting that the Lucas ratio framework and connections to modular forms established for the quadratic case may generalize, though deriving explicit minimum distributions for higher-order systems remains unexplored.
Beyond these generalizations, the interplay between different recurrence families appears to encode deep structural information. Preliminary computation reveals systematic patterns in how Fibonacci periods distribute across cyclotomic recurrences: for instance, the period $011$ appears modulo 2 in all $\Phi_{n}$ with $3\mid n$, and modulo 3 in all $\Phi_{3^{j}}$ for $j\geq 2$, while periods from class B2 primes (OEIS A053032, $p\geq 11$) rarely or never appear in cyclotomic landscapes. Whether this reflects deeper structural incompatibility between exponentially growing sequences and unit-root periodicity remains open. Studying complete period landscapes across polynomial families could reveal universal principles governing recurrence periodicity in finite rings. Fekete polynomials, which relate to cyclotomic polynomials and L-functions, as shown in [7], represent a particularly promising candidate for such investigations.
Finally, a central mystery remains: the classification parameter $\alpha$ that determines a prime’s class and governs the exact period count. While we established that the classification is intimately connected to the Legendre symbol $\left(\frac{5}{p}\right)$, with finer structure emerging via congruences modulo 20, the precise value of $\alpha$ for a given prime remains unknown. Understanding whether $\alpha$ relates to higher-order residue properties, Viswanath’s random recurrence through spectral properties, or other arithmetic invariants could eventually lead to an explicit description of the Pisano periods—resolving one of the longstanding open problems in the theory of Fibonacci sequences.